# Range-Limited Heaps’ Law for Functional DNA Words in the Human Genome

### v1

@@ Abstract
Heaps’ or Herdan’s law is a linguistic law describing the relationship between the vocabulary/dictionary size (type) and word counts (token) to be a power-law function. Its existence in genomes with certain definition of DNA words is unclear partly because the dictionary size in genome could be much smaller than that in a human language. We define a DNA word in a genome as a DNA coding region that codes for a protein domain. Using human chromosomes and chromosome arms as individual samples, we establish the existence of Heaps’ law in the human genome within limited range. Our definition of words in a genomic or proteomic context is different from that in large language models for DNA or protein sequences where words are usually short. Although an approximate power-law distribution of protein domain sizes due to gene duplication and the related Zipf’s law is well known, their translation to the Heaps’ law in DNA words is not automatic. Several other animal genomes are shown herein also to exhibit range-limited Heaps’ law with our definition of DNA words, though with various exponents, partially depending on their level of complexity. Investigation of Heaps’ law and its exponent value could provide an alternative narrative of reusage and redundancy of protein domains as well as creation of new protein domains from a linguistic perspective.
Abbreviations: CRCPD: coding regions that code for protein domains.
@@ 1 Introduction
In this study, we focus on another linguistic law, the Heaps law (Heaps, 1978) (or Herdan-Heaps law (Egghe, 2007)), which states that the number of unique words (vocabulary/dictionary size) as a function of text length is a power-law. Heaps’ law can be understood as a type(size)-token(size) (Herdan, 1960; Wetzel, 2009) relationship: type being a unique word, token being the usage of a word, i.e., the total number of types is the vocabulary size, and total number of tokens is the vocabulary usage. Heaps’ law can also be understood as a law of diminishing return. If one searches longer and longer texts for the purpose of finding new words, Heaps’ law indicates that the chance of finding new words becomes smaller and smaller as more texts are looked over. The latter meaning in Heaps law has been implied in the number of genes or gene-families discovered by sequencing more and more species’ genomes in a clade (Medini et al., 2020), or the number of single-nucleotide-polymorphisms (SNPs) discovered by sequencing more samples (Ionita-Laza et al., 2009).
@@
In order to apply a linguistic law to genomics, one first needs to find an entity corresponding to a word. A common practice is to consider (overlapping) k-mers with a prefixed k value as “words” due to its easiness to calculate, its use in alignment-free sequence comparison (Zielezinski et al., 2017), and its importance in many bioinformatics algorithms (Chikhi and Medvedev, 2013; Koren et al., 2017; Rahman et al., 2018). However, the problem with k-mers as words is that k-mers lack functional meaning, although there are attempts to filter k-mers by other criteria and by relaxing the fixed k value to obtain some meaning (Bussemaker et al., 2000; Moghaddasi et al., 2017). In the framework of semiotics (Hoopes, 2011), word is part of the signifier-signified pair that represents a meaning. Without a biological meaning, a proposed DNA or protein word would be much less interesting if we continue to discuss linguistic laws. Although the in-frame non-overlapping 3-mers within exons (codons) do have a biological meaning, k-mers, generally speaking, do not.
@@
Recently, large language models (Brant et al., 2007), such as bidirectional encoder representations from transformers (BERT) (Devlin et al., 2018) or generative pre-trained transformer (GPT) (Radford et al., 2019), have been applied to biomolecular sequences, in particular protein sequences (Rao et al., 2020; Madani et al., 2023). One aspect of natural language process is tokenization which means to delineate a character string by boundaries (Webster and Kit, 1992). For protein sequences, the tokenization is a process to find “words” (Ofer et al., 2021). The protein “words” found by this method are usually short, with average size of only a few (e.g. four) amino acids (Ferruz et al., 2022).
@@
In this study, we use DNA segments that code a protein domain (Wang et al., 2021) as our DNA “words”. Treating protein domains as protein words has been proposed before (Searls, 2002; Gimona, 2006; Scaiewicz and Levitt, 2015; Yu et al., 2019; Buchan and Jones, 2020). Proposals for DNA words were usually limited to k-mers (Brendel et al., 1986) which according to our discussion above is not ideal. There were studies that consider gene internal structure in DNA sequences as “grammar” or rules (Done and Searls, 1994), but the emphasis of these were not on the definition of words. We prefer protein domain reverse-transcribed to DNA sequences as DNA words. By doing so, our definition of DNA words has a biological meaning, i.e. coding for a protein domain.
@@
To investigate if Heaps’ law holds in the human genome, our task is to find all Coding Regions that Code for a Protein Domain (CRCPD). For this we need to count the number of times any CRCPD appears (number of tokens) and the number of different types of CRCPD appear (number of types). For one chromosome of the human genome, this pair of counts (#token, #type) will only give us one point. To see a relationship between token and type, we repeat the same counting process for each one of the 23 chromosomes and/or each one of the chromosome arms, so more points are produced. This may be chosen if more points are desired, as chromosomal arms may be seen as relatively independent entities, and different from the whole chromosomes where they belong.
@@ 2 Data and Methods
which after a logarithmic transformation on both side ($y^{\prime}=log(y)$, $x^{\prime}=log(x)$, $C^{\prime}=log(C)$): $y^{\prime}=C^{\prime}+\alpha x^{\prime}$ becomes a linear relationship. The exponent $\alpha$ is determined by a linear regression in R (the function lm). The AIC (for model selection) (Akaike, 1974; Li and Nyholt, 2001) calculation is carried out by the AICcmodavg package (cran.r-project.org/web/packages/AICcmodavg/), using the function aictab.
@@ 3.1 Heaps’ law for Pfam CRCPD holds true
The number of Pfam domain appearance in each autosomal and X chromosome in the human genome (version hg38) is obtained from the Pfam-in-GENECODE track of the UCSC genome browser. This would produce 23 points for checking the vocabulary size (type) and vocabulary usage (token) relationship and a regression. In order to increase the sample size, we also consider data from the p- and q-arm in a chromosome, as partitioned by the centromere for each metacentric (non-acrocentric) chromosome (41 points).
Fig.1(A) shows the number of types (unique number of Pfam domains) as a function of number of tokens (Pfam domain counts) for 23 human chromosomes in log-log scale. With the exception of one point, all other points scatter around a straightline, indicating a power-law and Heaps’ law. The outlier is chromosome 19, being lower in number of CRCPD types or higher in number of CRCPD tokens. The regression for log-token over log-type produces the slope (scaling exponent) $\alpha=0.658$ and $C=5.12$ in Eq.(1).
@@ 3.3 Chromosome 19 as outlier
The chromosome 19 has already been known to have the highest gene density compared to other chromosomes, more gene families, high GC content, higher density of repetitive sequences (Grimwood et al., 2004), extreme divergence with mouse genome (Castresana, 2002) but conserved within nonhuman primates (Harris et al., 2020).
We examine several chromosome level statistics to see which quantities make chromosome 19 an outlier. The chr19 has the highest GC content compared to other chromosomes, which might be related to the higher gene density, which itself is a product of widespread DNA/gene duplication.
If CRCPD statistics are examined directly, we find that the proportion of unique CRCPD (number of CRCPD types whose token, or copy number, is equal to 1) in chr19 does not stand out (see Table 1). The number of CRCPD types on chr19 is large, but not the largest. However, chr19 has the highest average copy number per CRCPD type, and also has the largest copy number for a CRCPD: zinc finger C2H2 domain appears 2805 times, which is an order of magnitude larger than the maximum copy number in the other chromosomes.
@@ 3.4 Extension to half of the chromosome arms
Figure 2 shows that with more data at the low end of the x-axis (the largest token count is 59 times the smallest token count), Heaps’ law remains to be true, with the fitting exponent $\alpha=0.75$. Removing the outliers due to chr.19, $\alpha=0.8$. Both of these slopes are larger than those in Fig.1.
@@
4 Results for Other Animal Genomes
Figure 3 shows the type-token scatter plot in log-log scale with points both representing a chromosome (black) or half of a chromosome (red). The practice of using half chromosomes is to increase the sample size and to increase the covering range of the x-axis.
To summarize, the results from these genomes where the protein domain track is available, Heaps’ law is generally true with exponent $\alpha<1$, despite potentially data quality issues such as the UniProt domain information, limited range of x-axis, and limited number of sample points. We also notice that when the low-end of the x-axis for the scatter plot is close to the origin (e.g. drosophila), the slope is relatively larger and closer to 1.
5 Artificial units generated by sampling tokens randomly and a systematic deviation from the power-law
The scatter plots in Fig.1 and Fig.2 are still too noisy, and the range of x-axis is still too limited. In order to have a clearer picture of the type-token relation in an even wider range, we cannot rely on the naturally provided units (“texts”) in chromosomes and chromosome arms, but to create artificial ones. We pool the appearance of CRCPD from all human chromosomes (62299 tokens). Then we randomly sample x number of tokens and check how many of them are unique (y number of types). A type-token plot from this random sampling is shown in Fig.4(A) in log-log scale.
If Heaps’ law holds exactly true, we should expect a straight line in the log-log plot in Fig.4. The linear regression (grey line) captures this trend (exponent 0.725 is less than one). However, a systematic deviation from the straight line, of a concave shape at both ends of the line, can be easily seen. When we add a quadratic term, it fits the data perfectly (with the coefficient for the quadratic term to be negative). The AIC of the quadratic regression, $-244.79$, is much better than that for the linear regression ($-93.08$), confirming the visual impression. This “log-log concavity” was previously reported (Bernhardsson et al., 2009; Lü et al., 2013; Font-Clos and Corral, 2015) and our DNA words result fully conforms it. For this reason, a better term for plots in Figs.1,2,3,4(A) could be “range-limited Heaps’ law”.

### v2

@@ Abstract
Heaps’ or Herdan’s law is a linguistic law describing the relationship between the vocabulary/dictionary size (type) and word counts (token) to be a power-law function. Its existence in genomes with certain definition of DNA words is unclear partly because the dictionary size in genome could be much smaller than that in a human language. We define a DNA word as a coding region in a genome that codes for a protein domain. Using human chromosomes and chromosome arms as individual samples, we establish the existence of Heaps’ law in the human genome within limited range. Our definition of words in a genomic or proteomic context is different from other definitions such as over-represented k-mers which are much shorter in length. Although an approximate power-law distribution of protein domain sizes due to gene duplication and the related Zipf’s law is well known, their translation to the Heaps’ law in DNA words is not automatic. Several other animal genomes are shown herein also to exhibit range-limited Heaps’ law with our definition of DNA words, though with various exponents. When tokens were randomly sampled and sample sizes reach to the maximum level, a deviation from the Heaps’ law was observed, but a quadratic regression in log-log type-token plot fits the data perfectly. Investigation of type-token plot and its regression coefficients could provide an alternative narrative of reusage and redundancy of protein domains as well as creation of new protein domains from a linguistic perspective.
Abbreviations: AIC: Akaike information criterion; CRCPD: coding regions that code for protein domains; GC content: percentage of guanine and cytosine bases.
@@ 1 Introduction
In this study, we focus on another linguistic law, the Heaps law (Heaps, 1978) (or Herdan-Heaps law (Egghe, 2007)), which states that the number of unique words (vocabulary/dictionary size) as a function of text length is a power-law. Heaps’ law can be understood as a (number of)type-(number of)token (Herdan, 1960; Wetzel, 2009) relationship: type being a unique word, token being usage of a word, i.e., the total number of types is the vocabulary size, and total number of tokens is the vocabulary usage. Heaps’ law can also be understood as a law of diminishing return (Petersen et al., 2012). If one searches longer and longer texts for the purpose of finding new words, Heaps’ law indicates that the chance of finding new words becomes smaller and smaller as more texts are looked over. The latter meaning in Heaps law has been implied in the number of genes or gene-families discovered by sequencing more and more species’ genomes in a clade (Medini et al., 2020), or the number of single-nucleotide-polymorphisms (SNPs) discovered by sequencing more samples (Ionita-Laza et al., 2009).
@@
In order to apply a linguistic law to genomics, one first needs to find an entity corresponding to a word. A common practice is to consider (overlapping) k-mers with a prefixed k value as “words” due to its easiness to calculate, its use in alignment-free sequence comparison (Zielezinski et al., 2017; Li et al., 2019), and its importance in many bioinformatics algorithms (Chikhi and Medvedev, 2013; Koren et al., 2017; Rahman et al., 2018). However, the problem with all fixed-length k-mers as words is that k-mers in general lack functional meaning.
Instead of all k-mers, attempts were made to select only the over-represented k-mers, based on the frequency expectation from (k-1)-mers, as words or motifs (Brendel et al., 1986; Phillips, et al., 1987; Vilo, 2002; Apostolico et al., 2003; Gatherer, 2007). There are other motif detection methods based on statistical mechanics (Bussemaker et al., 2000; Moghaddasi et al., 2017). However, these are postulated units only. In fact, non-coding regions tend to have relatively more over-represented oligonucleotides than coding regions (Frontali and Pizzi, 1999). Many of over-represented oligonucleotides in noncoding regions may be due to error accumulation (e.g. slippage errors) tolerated by purifying selection, but some others may have functional implications and thus deserve the name “words”, as the mirror-symmetrical words found in introns (Brendel et al., 1986).
In the framework of semiotics (Hoopes, 2011), word is part of the signifier-signified pair that represents a meaning. Without a biological meaning, a proposed DNA or protein word would be much less interesting if we continue to discuss linguistic laws. Although the in-frame non-overlapping 3-mers within exons (codons) do have a biological meaning (Mukhopadhyay et al., 2006), k-mers, even the over-represented ones, may or may not.
@@
Recently, large language models (Brant et al., 2007), such as bidirectional encoder representations from transformers (BERT) (Devlin et al., 2018) or generative pre-trained transformer (GPT) (Radford et al., 2019), have been applied to biomolecular sequences, in particular protein sequences (Rao et al., 2020; Madani et al., 2023; Nijkamp et al., 2023). One aspect of natural language process is tokenization which means to delineate a character string by boundaries (Webster and Kit, 1992). For protein sequences, the tokenization is a process to find “words” (Ofer et al., 2021). The protein “words” found by this method are usually short, with average size of only a few (e.g. four) amino acids (Ferruz et al., 2022; Dotan et al., 2024).
@@
In this study, we use DNA segments that code a protein domain (Wang et al., 2021) as our DNA “word tokens”. Treating protein domains as protein words has been proposed before (Searls, 2002; Gimona, 2006; Scaiewicz and Levitt, 2015; Yu et al., 2019; Buchan and Jones, 2020). There were studies that consider gene internal structure in DNA sequences as “grammar” or rules (Dong and Searls, 1994), but the emphasis of these were not on the definition of words. We prefer to define DNA words as substrings of gene sequences delimited by the boundaries of the protein domains within the proteins encoded by those genes. By doing so, our definition of DNA words has a biological meaning, i.e. coding for a protein domain.
@@
To investigate if Heaps’ law holds in the human genome, our task is to find all Coding Regions that Code for a Protein Domain (CRCPD). For this we need to, for any genomic region, count the number of all CRCPDs (number of tokens) and the number of distinct CRCPDs among them (number of types). For one chromosome of the human genome, this pair of counts (#token, #type) will only give us one point. To see a relationship between token and type, we repeat the same counting process for each one of the 23 chromosomes and/or each one of the chromosome arms, so more points are produced. This may be chosen if more points are desired, as chromosomal arms may be seen as relatively independent entities, and different from the whole chromosomes where they belong.
@@ 2 Data and Methods
which after a logarithmic transformation on both side ($y^{\prime}=log(y)$, $x^{\prime}=log(x)$, $C^{\prime}=log(C)$): $y^{\prime}=C^{\prime}+\alpha x^{\prime}$ becomes a linear relationship. The exponent $\alpha$ is determined by a linear regression in R (the function lm). For model selection the Akaike information criterion (AIC) (Akaike, 1974; Li and Nyholt, 2001) calculation is carried out by the AICcmodavg package (cran.r-project.org/web/packages/AICcmodavg/), using the function aictab. For quadratic regression, we fit the equation $y^{\prime}=C+\alpha x^{\prime}+\beta x^{\prime 2}$.
The following notation may gives readers an idea of the type of data we have. Suppose there are N tokens (appearance of CRCPDs: coding for Pfam domains), and K unique ones. $N=\sum_{i=1}^{K}P_{k}$, where $P_{k}$ is the multiplicity of word type-$k$. For any genomic region (chromosome, chromosome arm, or half-arm), we count the number of tokens $x_{i}$ ($i$ is the index for sample point) and $y_{i}$ of them are unique. When we have enough samplings $i=1,2,\cdots,I$, these ($x_{i}$, $y_{i}$) pairs can be log-transformed and used for a linear regression. We may not be constrained by genomic regions, but sample $x_{i}$ tokens genome-wide from the pool of N tokens (sampling without replacement) and among them $y_{i}$ ones are unique. This can be repeated I times with ever larger $x_{i}$ values. The same ($x_{i}$,$y_{i}$) pairs can be used to check Heaps’ law.
@@ 3.1 Heaps’ law for Pfam CRCPD holds true
The number of Pfam domain appearance, as our example of CRCPD, in each autosomal and X chromosome in the human genome (version hg38) is obtained from the Pfam-in-GENECODE track of the UCSC genome browser. This would produce 23 points for checking the vocabulary size (type) and vocabulary usage (token) relationship and a regression. In order to increase the sample size, we also consider data from the p- and q-arm in a chromosome, as partitioned by the centromere for each metacentric (non-acrocentric) chromosome (41 points).
Fig.1(A) shows the number of types (number of unique/distinct Pfam domains) as a function of number of tokens (Pfam domain counts) for 23 human chromosomes in log-log scale. With the exception of one point, all other points scatter around a straight-line, indicating a power-law and Heaps’ law. The outlier is chromosome 19, being lower in number of CRCPD types or higher in number of CRCPD tokens. The regression for log-token over log-type produces the slope (scaling exponent) $\alpha=0.658$ and $C=5.12$ in Eq.(1).
@@
Figure 1: Type-token plot, where $x$-axis is the number of CRCPD tokens (vocabulary usage) and $y$-axis is the number of unique CRCPD type (vocabulary size), for human genome. (A) Each point corresponds to a human chromosome, in log-log scale. (B) Each point corresponds to a human chromosome arm, in log-log scale. (C) Each point corresponds to either a human chromosome or a chromosome arm, in log-log scale. (D) Same as (C) but in linear-linear scale. Chromosome 19 is clearly indicated as an outlier in all panels.
@@ 3.2 The protein domains that are mostly responsible for the redundancy
ch most_freq CN 2nd CN 3rd CN num_type per(CN=1) mean(CN) max(CN)
1 Olduvai 215 zf-C2H2 214 Sushi 186 1759 64.7% 3.597 215
2 I-set 224 fn3 156 Ig_3 136 1355 63.8% 3.373 224
3 zf-C2H2 217 WD40 72 Ig_3 70 1178 65.3% 2.93 217
4 Cadherin 101 zf-C2H2 92 Ank 50 857 67.1% 2.797 101
5 Cadherin 334 zf-C2H2 134 Cadherin_2 55 951 68.2% 3.121 334
6 zf-C2H2 162 fn3 62 zf-C2H2_6 57 1057 65.8% 2.807 162
7 zf-C2H2 324 zf-C2H2_6 91 V-set 58 957 66.1% 2.883 324
8 zf-C2H2 171 Sushi 56 Ank 52 787 67.9% 2.663 171
9 zf-C2H2 144 I-set 57 Ig_3 57 856 67.4% 2.974 144
10 zf-C2H2 103 Ank 72 Ank_3 44 854 66% 2.56 103
11 7tm_1 196 7tm_4 180 I-set 70 1136 65.8% 3.109 196
12 zf-C2H2 127 WD40 48 Ank 40 1100 65.8% 2.576 127
13 Collagen 28 Cadherin 21 zf-C2H2 19 475 69.7% 2.027 28
14 V-set 96 WD40 41 7tm_1 40 736 67.1% 2.455 96
15 zf-C2H2 57 EGF_CA 46 hEGF 41 727 68.1% 2.611 57
16 zf-C2H2 292 zf-C2H2_6 71 WD40 44 914 67.2% 2.534 292
17 zf-C2H2 107 Keratin_B2_2 73 WD40 42 1151 65.2% 2.487 107
18 zf-C2H2 75 Cadherin 45 Laminin_EGF 30 370 69.5% 2.473 75
19 zf-C2H2 2805 zf-C2H2_6 910 KRAB 228 1114 65.4% 6.684 2805
20 zf-C2H2 107 zf-C2H2_6 25 Laminin_EGF 21 674 72.6% 2.068 107
21 Keratin_B2_2 47 I-set 17 ig 16 291 69.4% 2.131 47
22 V-set 39 zf-C2H2 27 EGF_CA 16 590 68.1% 2.171 39
X zf-C2H2 136 Collagen 32 MAGE 32 781 59.3% 2.77 136
all zf-C2H2 5493 zf-C2H2_6 1609 I-set 900 6839 44.3% 9.109 5483
Table 1: Some human chromosome level statistics on protein domains (CN: copy number, or number of CRCPD token for a specific CRCPD type): the protein domain with the highest, second highest, and the third highest, CN; number of unique CRCPDs; proportion of CRCPD types that have CN=1 (which could be called a singleton); average copy number (or token-over-type ratio); and maximum copy number. The maximum (minimum) value in each column is marked with bold (italics).
@@ 3.3 Chromosome 19 as outlier
Chromosome 19 (Chr19) has already been known to have the highest gene density compared to other chromosomes, more gene families, high GC content, higher density of repetitive sequences (Grimwood et al., 2004), extreme divergence with mouse genome (Castresana, 2002) but conserved within nonhuman primates (Harris et al., 2020).
We examine several chromosome level statistics to see which quantities make chromosome 19 an outlier. Chr19 has the highest GC content compared to other chromosomes, which might be related to the higher gene density, which itself is a product of widespread DNA/gene duplication.
If CRCPD statistics are examined directly, we find that the proportion of unique CRCPD in Chr19 does not stand out (see Table 1). The number of CRCPD types on Chr19 is large, but not the largest. However, Chr19 has the highest average copy number per CRCPD, and also has the largest copy number for a CRCPD: zinc finger C2H2 domain appears 2805 times, which is an order of magnitude larger than the maximum copy number in the other chromosomes.
@@ 3.4 Extension to half of the chromosome arms
Figure 2 shows that with more data at the low end of the x-axis (the largest token count is 59 times the smallest token count), Heaps’ law remains to be true, with the fitting exponent $\alpha=0.75$. Removing the outliers due to Chr19, $\alpha=0.8$. Both of these slopes are larger than those in Fig.1.
@@
Figure 2: An extension of Fig.1(C). Besides human chromosomes and chromosome arms, half-chromosome-arms (blue) are used to count CRCPD tokens and types.
4 Robustness of the results
4.1 Random shuffle of CRCPDs
In order to check if the sampling units of chromosomes, chromosome arms, or half arms, play any role in the observed power-law trend in Figs.1,2, we carried out two experiments. The first is that we shuffle the CRCPDs (tokens) genome-wide once, while keeping the chromosome location intact. Then we repeat the same CRCPD-type count in the same chromosomes, chromosome arms, and half arms. In the next subsection, we will randomly sample CRCPDs without replacement. Then the CRCPD-type count is plotted against the number of CRCPD token sampled.
Fig.3 shows the type-token relationship when CRCPDs are shuffled in the whole genome. The black dots are identical to those in Fig.2, some dots represent chromosomes, others represent chromosome arms or half arms. The red dots are the similar type counts in the same genomic segments (chromosomes, arms, half arms) when the shuffling has taken place.
It can be seen from Fig.3 that after shuffling, each genomic unit tends to have more CRCPD types than the original data (red dots being above black dots). It means that CRCPD redundancy tends to be locally clustered. However, the shuffling of CRCPDs does not change the fact that type and token are in a power-law relationship. The regression coefficient is 0.747 for the shuffled data, very close to the original values of 0.747 (or 0.8 when Chr19 is removed) for the original data.
Figure 3: CRCPD types count in chromosomes, chromosome arms, and half arms after CRCPD tokens are shuffled genome-wide (red dots), as a function of CRCPD token counts in these genomic units. The black dots are idential to those in Fig.2, for the same type-token counts when CRCPDs are not shuffled.
4.2 Random sampling from CRCPD pool and a systematic deviation from the power-law
In the second approach to test the robustness of our result, we create artificial sampling units by randomly picking CRCPD tokens from the genome. This new way of sampling tokens would free us from the restriction of chromosomes or any genomic regions. Also, the range of x-axis in type-token plot can be greatly expanded. We pool CRCPDs from all human chromosomes (62299 tokens). Then we randomly sample x number of tokens and check how many of them are unique (y number of types). A type-token plot from this random sampling is shown in Fig.4(A) in log-log scale. Note that the x-range in Fig.4(A) is much more expanded compared to Figs.1,2.
If Heaps’ law holds exactly true, we should expect a straight line in the log-log plot in Fig.4. The linear regression (grey line) captures this trend (exponent 0.725 is less than one). However, a systematic deviation from the straight line, of a concave shape at both ends of the line, can be easily seen. When we add a quadratic term, it fits the data perfectly (with the coefficient for the quadratic term to be negative). The AIC of the quadratic regression, $-244.79$, is much better than that for the linear regression ($-93.08$), confirming the visual impression. This “log-log concavity” was previously reported (Bernhardsson et al., 2009; Lü et al., 2013; Font-Clos and Corral, 2015) and our DNA words result fully conforms it. For this reason, a better term for plots in Figs.1,2,6,4(A) could be “range-limited Heaps’ law”.